# The Identification Power of

### v1

@@ 2.2 Distributional Treatment Effect Parameters
$\theta_{o}\equiv\mathbb{E}\left[\psi(Y_{1},Y_{0})\right],$
where $\psi:\mathbb{R}^{2}\to\mathbb{R}$ is a specified function. With various specification of $\psi$, this formulation encompasses a wide range of DTE parameters, as illustrated below:
@@
When $\psi(y_{1},y_{0})=\mathbf{1}\{y_{1}>y_{0}\}$, we have $\theta_{o}=\mathbb{P}(Y_{1}>Y_{0})$, the fraction of individuals who benefit from the treatment. This parameter captures the distributional impact of the treatment. Importantly, even when the ATE $\mathbb{E}[Y_{1}-Y_{0}]$ is positive, $\mathbb{P}(Y_{1}>Y_{0})$ may be small, indicating that few individuals experience large gains while many incur losses. Focusing on $\mathbb{P}(Y_{1}>Y_{0})$ is thus essential to avoid interventions that are unfavorable to the majority.
@@
We can also consider the ATE for disadvantaged individuals, defined as $\mathbb{E}[Y_{1}-Y_{0}\mid Y_{0}\leq c]$ for a fixed threshold $c$. This can be expressed by the function $\psi(y_{1},y_{0})=(y_{1}-y_{0})\cdot\mathbf{1}\{y_{0}\leq c\}/\mathbb{P}(Y_{0}\leq c)$. The subpopulation with $Y_{0}\leq c$ represents individuals whose baseline outcomes fall below the disadvantage threshold $c$ (e.g., a poverty line). The nuisance parameter $\mathbb{P}(Y_{0}\leq c)$ appearing in $\psi(y_{1},y_{0})$ is point-identified from experimental data under the maintained assumptions as $\mathbb{P}(Y_{0}\leq c)=\mathbb{P}(Y\leq c\mid D=0,G=\textit{exp})$.
@@
When $\psi$ is specified as $\psi(y_{1},y_{0})=\mathbf{1}\{y_{1}>c,y_{0}\leq c\}/\mathbb{P}(Y_{0}\leq c)$ for a fixed threshold $c$, the target parameter $\theta_{o}$ becomes $\theta_{o}=\mathbb{P}(Y_{1}>c\mid Y_{0}\leq c)$, which represents the probability of upward mobility among individuals initially below $c$. This parameter captures the likelihood that individuals whose baseline outcomes fall below the threshold—such as a poverty line—will exceed it under treatment. Unlike the ATE, this measure focuses specifically on distributional gains among the disadvantaged.
Example 2.5 (Correlation Between the Two Potential Outcomes).
@@
When $\psi(y_{1},y_{0})=(y_{1}-\mathbb{E}[Y_{1}])(y_{0}-\mathbb{E}[Y_{0}])/\sqrt{\text{Var}(Y_{1})\text{Var}(Y_{0})}$, the target parameter $\theta_{o}$ corresponds to $\text{Cor}(Y_{1},Y_{0})$, the correlation between the two potential outcomes. This parameter reflects the degree of similarity in individual-level responses to the two interventions, offering insight into treatment effect heterogeneity. The nuisance components $\mathbb{E}[Y_{d}]$ and $\text{Var}(Y_{d})$ for each $d\in\{0,1\}$ are point-identified from experimental data under the maintained assumptions.
@@
Note that none of the DTE parameters introduced above can be point-identified, even with experimental data. This is because their identification generally requires knowledge of the joint distribution of the two potential outcomes, $F_{Y_{1}Y_{0}}^{\ast}$, whereas the experimental data only identify the marginal distributions, $(F_{Y_{1}}^{\ast},F_{Y_{0}}^{\ast})$.
@@
We can also consider DTE parameters for various subpopulations: the parameter for each self-section group, $\theta_{o,s}\equiv\mathbb{E}[\psi(Y_{1},Y_{0})\mid S=s]$; the parameter conditional on covariates, $\theta_{o,x}\equiv\mathbb{E}[\psi(Y_{1},Y_{0})\mid X=x]$; and the parameter for each data source population, $\theta_{o,g}\equiv\mathbb{E}[\psi(Y_{1},Y_{0})\mid G=g]$ for $g\in\{\mathrm{exp},\mathrm{obs}\}$. In particular, $\theta_{o,s}$ captures the treatment effect for individuals who self-select into treatment ($s=1$) or no-treatment ($s=0$), and is relevant in many empirical applications.
These subpopulation parameters can be incorporated into our framework with slight modifications. For example, given a fixed $\psi$, the parameter $\theta_{o,g}$ for each data source population can be expressed as $\theta_{o,g}=\mathbb{E}\left[\psi(Y_{1},Y_{0})\cdot\mathbf{1}\{G=g\}/\mathbb{P}(G=g)\right]$, where $\mathbb{P}(G=g)$ is point-identified from the observed data. By defining $\tilde{\psi}(y_{1},y_{0})=\psi(y_{1},y_{0})\cdot\mathbf{1}\{G=g\}/\mathbb{P}(G=g)$, the parameter $\theta_{o,g}$ takes the form $\mathbb{E}\left[\tilde{\psi}(Y_{1},Y_{0})\right]$, and can therefore be treated analogously to $\theta_{o}$ in our analysis. As for $\theta_{o,s}$, Lemma C.1 in the supplementary appendix shows that $\theta_{o,s}=\mathbb{E}[\psi(Y_{1},Y_{0})|D=s,G=\mathrm{obs}]$, which can be equivalently written as $\theta_{o,s}=\mathbb{E}[\psi(Y_{1},Y_{0})\cdot\mathbf{1}\{D=s,G=\mathrm{obs}\}/\mathbb{P}(D=s,G=\mathrm{obs})]$. Hence, $\theta_{o,s}$ can likewise be handled in the same manner as $\theta_{o}$.
@@
In what follows, we examine the partial identification of $\theta_{o}$ using experimental data alone and using both experimental and observational data.
@@ 2.3 Identified Sets for DTE parameters under Experimental and Combined Data
We begin by formally defining the identified sets for the DTE parameter $\theta_{o}$ under two data scenarios: experimental data only and the combined experimental and observational data. We first consider the case based solely on experimental data.
Let $\mathcal{F}^{\dagger}$ denote the class of CDFs for all defined variables $(Y_{1},Y_{0},Y,D,S,G,X)$ that satisfy Assumptions 2.12.5.444Formally, $\mathcal{F}^{\dagger}$ is the set of all CDFs $F$ of $(Y_{1},Y_{0},Y,D,S,G,X)$ that satisfy Assumptions 2.12.5 with $F^{\ast}$ replaced by $F$. We define the identified set for the joint CDF $F^{\ast}$ under the experimental data as
@@
where $F_{YDX|G}^{\ast}(\cdot,\cdot,\cdot|\mathrm{exp})$ is the distribution of the observables $(Y,D,X)$ in the experimental data, and $F_{X}^{\ast}$ is the marginal distribution of covariates in the overall population. The set $\mathcal{F}_{\mathrm{exp}}^{\ast}$ comprises all CDFs that satisfy the maintained assumptions and are consistent with the distribution of the observed experimental data and the marginal distribution of covariates in the population.
We assume that $F_{X}^{\ast}$ is known, as it is required to adjust for potential imbalances in the covariate distribution between the experimental and observational data and to facilitate meaningful comparisons between the identified sets. In practice, such covariate information is often available from external sources, such as demographic data for the overall population. There are also some cases where knowledge of $F_{X}^{\ast}$ is not required—for example, when the covariate distribution is balanced between the data sources (i.e., $F_{X|G}^{\ast}(\cdot\mid\mathrm{exp})=F_{X|G}^{\ast}(\cdot\mid\mathrm{obs})$), or when the parameter of interest is the covariate-conditional effect $\theta_{o,x}$ or the effect $\theta_{o,\mathrm{exp}}$ defined for the experimental population. In such cases, $F_{X}^{\ast}$ is either identified from the experimental data or irrelevant to the target parameter.
@@
This set comprises all parameter values that are attainable under some distribution in $\mathcal{F}_{\mathrm{exp}}^{\ast}$. Any parameter value outside $\Theta_{I}$ is incompatible with either the experimental data or the maintained assumptions.
We next consider the case where both experimental and observational data are available. As before, we begin by defining the identified set for the joint CDF $F^{\ast}$ under the combined data, in analogy to $\mathcal{F}_{\mathrm{exp}}^{\ast}$:
@@
where $F_{YDGX}^{\ast}$ is the joint distribution of the observables $(Y,D,G,X)$ in the combined data. This set consists of all CDFs that satisfy the maintained assumptions and are consistent with the observed distribution $F_{YDGX}^{\ast}$. Note that $\mathcal{F}^{\ast}\subseteq\mathcal{F}_{\mathrm{exp}}^{\ast}$, since the condition $F_{YDGX}=F_{YDGX}^{\ast}$ implies both $F_{YDX|G}(\cdot,\cdot,\cdot\mid\mathrm{exp})=F_{YDX|G}^{\ast}(\cdot,\cdot,\cdot\mid\mathrm{exp})$ and $F_{X}=F_{X}^{\ast}$.
@@
This set comprises all parameter values that are attainable under some CDF in $\mathcal{F}^{\ast}$. Any value outside $\Theta_{IC}$ contradicts either the maintained assumptions or the combined data.
Since $\mathcal{F}^{\ast}\subseteq\mathcal{F}_{\mathrm{exp}}^{\ast}$, we have $\Theta_{IC}\subseteq\Theta_{I}$; that is, the identified set under the combined data is no larger than that under experimental data alone. Our primary interest, however, lies in whether the strict inclusion $\Theta_{IC}\subset\Theta_{I}$ holds—that is, whether combining experimental and observational data strictly tightens the identified set for $\theta_{o}$. This is a nontrivial question, as a strict inclusion $\mathcal{F}^{\ast}\subset\mathcal{F}_{\mathrm{exp}}^{\ast}$ at the level of CDFs does not necessarily imply a strict inclusion $\Theta_{IC}\subset\Theta_{I}$ at the level of the target parameter. We examine this question in the following sections.
@@ 3 Characterization of the Identified Sets
To investigate whether the strict inclusion $\Theta_{IC}\subset\Theta_{I}$ holds, the definitions of $\Theta_{I}$ and $\Theta_{IC}$ given in equations (1) and (2) are too abstract to draw any meaningful insight. We therefore seek to characterize these identified sets in a more interpretable form. To this end, we make use of the concept of bivariate copulas (Sklar, 1959), which will serve as a key tool in this and the following sections.
@@ 3.1 Characterization of $\Theta_{I}$ for Experimental Data
We begin with the case of experimental data. From the randomized treatment assignment and Assumption 2.4 (external validity), we can identify the conditional marginal CDFs of the potential outcomes given covariates, $(F_{Y_{1}|X}^{\ast},F_{Y_{0}|X}^{\ast})$, as $F_{Y_{d}|X}^{\ast}(\cdot|x)=F_{Y|DGX}^{\ast}(\cdot|d,\mathrm{exp},x)$ for d=0,1d=0,1.555This follows from the equalities $F_{Y_{d}|X}^{\ast}(\cdot|x)=F_{Y_{d}|GX}^{\ast}(\cdot|\mathrm{exp},x)=F_{Y|DGX}^{\ast}(\cdot|d,\mathrm{exp},x)$, where the first and second equalities follow from Assumptions 2.4 and 2.1, respectively. Since the marginal distribution of covariates $F_{X}^{\ast}$ is assumed to be known, we can identify the joint distribution of each potential outcome and covariates, $(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast})$.
Let $\mathcal{C}$ denote the class of all bivariate copula functions. For each $x\in\mathcal{X}$, let $C^{\ast}(\cdot,\cdot|x)\in\mathcal{C}$ denote the true conditional copula given $x$ that reproduces the true conditional joint distribution $F_{Y_{1}Y_{0}|X}^{\ast}(y_{1},y_{0}|x)$ from the marginals $(F_{Y_{1}|X}^{\ast}(y_{1}|x),F_{Y_{0}|X}^{\ast}(y_{0}|x))$:
@@
where the existence of such a copula function is guaranteed by Sklar’s theorem (e.g., Nelsen, 2006, Theorem 2.3.3). [^6] Sklar’s theorem (Sklar, 1959) states that any joint distribution $F_{Y_{1}Y_{0}}$ can be expressed as a copula function of its marginals: $F_{Y_{1}Y_{0}}(y_{1},y_{0})=C(F_{Y_{1}}(y_{1}),F_{Y_{0}}(y_{0}))$, for some copula function $C$. Conversely, given any marginal distributions $F_{Y_{1}}$ and $F_{Y_{0}}$, the function $C(F_{Y_{1}}(y_{1}),F_{Y_{0}}(y_{0}))$, for any copula $C$, defines a valid bivariate distribution with marginals $F_{Y_{1}}$ and $F_{Y_{0}}$. Sklar’s theorem plays an important role in partial identification analysis. Using the true conditional copula $C^{\ast}(\cdot,\cdot|x)$, the true parameter $\theta_{o}$ can be expressed as
@@
The true copula function $C^{\ast}(\cdot,\cdot|x)$ is unknown. However, by replacing $C^{\ast}(\cdot,\cdot|x)$ with all possible copula functions $C(\cdot,\cdot|x)\in\mathcal{C}$, we obtain the identified set for $\theta_{o}$ based on the knowledge of $(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast})$ as follows:
@@
where, as implied by Sklar’s theorem, the set $\bigl{\{}C\bigl{(}F_{Y_{1}|X}^{\ast}(\cdot|x),F_{Y_{0}|X}^{\ast}(\cdot|x)\big{|}x\bigr{)}:C(\cdot,\cdot|x)\in\mathcal{C}\bigr{\}}$ corresponds to the collection of conditional joint CDFs of $(Y_{1},Y_{0})$ given $X=x$ that share the marginal CDFs $\bigl{(}F_{Y_{1}|X}^{\ast}(\cdot|x),F_{Y_{0}|X}^{\ast}(\cdot|x)\bigr{)}$. Fan et al. (2017) show that the covariates $X$ themselves contribute to shrinking the identified set by restricting the admissible dependence structures.
The following proposition formally establishes that the identified set $\widetilde{\Theta}_{I}$, defined based on the joint distributions $(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast})$, coincides with the identified set $\Theta_{I}$ under the experimental data.
@@
The identified set $\Theta_{I}$ under the experimental data is equivalent to the identified set $\widetilde{\Theta}_{I}$ based on the knowledge of $(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast})$: that is, $\Theta_{I}=\widetilde{\Theta}_{I}$.
While this result would not be surprising, it formally shows that the identified set $\Theta_{I}$ is fully characterized by the joint distributions $(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast})$. [^7] While many studies refer to $\widetilde{\Theta}_{I}$ as the identified set under the experimental data, we formally show that $\widetilde{\Theta}_{I}$— defined based on $(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast})$— indeed coincides with the identified set $\Theta_{I}$ defined based on the experimental data.

### v4

@@ 2.2 Distributional Treatment Effect Parameters
$\theta_{o}\equiv\mathbb{E}\left[\psi(Y_{1},Y_{0})\right],$ (1)
where $\psi:\mathbb{R}^{2}\to\mathbb{R}$ is a given function. With various specifications of $\psi$, this formulation encompasses a wide range of DTE parameters, as illustrated below.
@@
When $\psi(y_{1},y_{0})=\mathbf{1}\{y_{1}>y_{0}\}$, we have $\theta_{o}=\mathbb{P}(Y_{1}>Y_{0})$, the fraction of individuals who benefit from the treatment. This parameter captures the distributional impact of the treatment. Importantly, even if the ATE $\mathbb{E}[Y_{1}-Y_{0}]$ is positive, the fraction $\mathbb{P}(Y_{1}>Y_{0})$ may be small, indicating that a small number of individuals experience large gains while many experience no gains or losses. Focusing on $\mathbb{P}(Y_{1}>Y_{0})$ is therefore essential to avoid interventions that are unfavorable for the majority.
@@
Consider the ATE for disadvantaged individuals, defined as $\mathbb{E}[Y_{1}-Y_{0}\mid Y_{0}\leq c]$ for a fixed threshold $c$. This parameter can be expressed in the form of equation (1) with $\psi(y_{1},y_{0})=(y_{1}-y_{0})\cdot\mathbf{1}\{y_{0}\leq c\}/\mathbb{P}(Y_{0}\leq c)$. The subpopulation with $Y_{0}\leq c$ represents individuals whose outcomes fall below the disadvantage threshold $c$ (e.g., a poverty line) without treatment. The nuisance parameter $\mathbb{P}(Y_{0}\leq c)$ in $\psi(y_{1},y_{0})$ is point-identified from the experimental data under the maintained assumptions as $\mathbb{P}(Y_{0}\leq c)=\mathbb{P}(Y\leq c\mid D=0,G=\textit{exp})$.
@@
When $\psi(y_{1},y_{0})=\mathbf{1}\{y_{1}>c,y_{0}\leq c\}/\mathbb{P}(Y_{0}\leq c)$ for a fixed threshold $c$, the target parameter $\theta_{o}$ becomes $\theta_{o}=\mathbb{P}(Y_{1}>c\mid Y_{0}\leq c)$. This parameter represents the probability of upward mobility among individuals whose outcomes would fall below $c$ in the absence of treatment. Specifically, it captures the likelihood that individuals whose baseline outcomes are below a given threshold—such as a poverty line—will exceed that threshold when treated. Unlike the ATE, this parameter focuses on distributional gains among the disadvantaged subpopulation.
Example 2.5 (Correlation between Two Potential Outcomes).
@@
When $\psi(y_{1},y_{0})=(y_{1}-\mathbb{E}[Y_{1}])(y_{0}-\mathbb{E}[Y_{0}])/\sqrt{\text{Var}(Y_{1})\text{Var}(Y_{0})}$, we have $\theta_{o}=\text{Cor}(Y_{1},Y_{0})$, the correlation between the two potential outcomes. This parameter reflects the degree of similarity in individual-level responses to the two interventions, offering insights into treatment effect heterogeneity. The nuisance components $\mathbb{E}[Y_{d}]$ and $\text{Var}(Y_{d})$ for $d\in\{0,1\}$ are point-identified from the experimental data under the maintained assumptions.
See also Fan et al. (2017) for additional examples.
@@
Note that none of the DTE parameters introduced above can be point-identified even with experimental data. Their identification generally requires knowledge of the joint distribution of the two potential outcomes, $F_{Y_{1}Y_{0}}^{\ast}$, whereas experimental data identify only the marginal distributions, $(F_{Y_{1}}^{\ast},F_{Y_{0}}^{\ast})$.
@@
We can also define DTE parameters for various subpopulations: by self-selection type, $\theta_{o,s}\equiv\mathbb{E}[\psi(Y_{1},Y_{0})\mid S=s]$; conditional on covariates, $\theta_{o,x}\equiv\mathbb{E}[\psi(Y_{1},Y_{0})\mid X=x]$; and by data source, $\theta_{o,g}\equiv\mathbb{E}[\psi(Y_{1},Y_{0})\mid G=g]$ for $g\in\{\mathrm{exp},\mathrm{obs}\}$. In particular, $\theta_{o,s}$ captures the treatment effect for individuals who self-select into treatment ($s=1$) or no-treatment ($s=0$), a quantity that is relevant in many empirical applications.
These subpopulation parameters can be incorporated into our framework with slight modifications. For example, the parameter $\theta_{o,g}$ can be expressed as
$\theta_{o,g}=\mathbb{E}\left[\psi(Y_{1},Y_{0})\cdot\frac{\mathbf{1}\{G=g\}}{\mathbb{P}(G=g)}\right],$
where $\mathbb{P}(G=g)$ is point-identified from the observed data. Defining $\tilde{\psi}(y_{1},y_{0})=\psi(y_{1},y_{0})\cdot\mathbf{1}\{G=g\}/\mathbb{P}(G=g)$, we can write $\theta_{o,g}=\mathbb{E}[\tilde{\psi}(Y_{1},Y_{0})]$, so it can be analyzed analogously to $\theta_{o}$. As for $\theta_{o,s}$, Lemma C.1 in the appendix shows that
$\theta_{o,s}=\mathbb{E}[\psi(Y_{1},Y_{0})|D=s,G=\mathrm{obs}]=\mathbb{E}\left[\psi(Y_{1},Y_{0})\cdot\frac{\mathbf{1}\{D=s,G=\mathrm{obs}\}}{\mathbb{P}(D=s,G=\mathrm{obs})}\right].$
Hence, $\theta_{o,s}$ can also be handled in the same manner as $\theta_{o}$.
@@
In what follows, we examine the partial identification of $\theta_{o}$ based on experimental data alone and based on the combination of experimental and observational data.
@@ 2.3 Identified Sets for DTE Parameters under Experimental and Combined Data
We formally define the identified sets for the DTE parameter $\theta_{o}$ under two data scenarios: (i) experimental data alone and (ii) the combination of experimental and observational data. We begin with the case of experimental data only.
Let $\mathcal{F}^{\dagger}$ denote the class of CDFs for all defined variables $(Y_{1},Y_{0},Y,D,S,G,X)$ that satisfy Assumptions 2.12.5.333Formally, $\mathcal{F}^{\dagger}$ is the set of all CDFs $F$ of $(Y_{1},Y_{0},Y,D,S,G,X)$ that satisfy Assumptions 2.12.5 with $F^{\ast}$ replaced by $F$. We begin by defining the identified set for the joint CDF $F^{\ast}$ under the experimental data as
@@
where $F_{YDX|G}^{\ast}(\cdot,\cdot,\cdot|\mathrm{exp})$ is the distribution of the observables $(Y,D,X)$ in the experimental data, and $F_{X}^{\ast}$ is the marginal distribution of $X$ in the entire population. Thus, $\mathcal{F}_{\mathrm{exp}}^{\ast}$ consists of all CDFs that satisfy the maintained assumptions and are consistent with the distribution of the observed experimental data and the marginal distribution of $X$ in the entire population.
We assume that $F_{X}^{\ast}$ is known, as this adjustment is necessary to account for potential imbalances in the covariate distributions between the experimental and observational data and to enable meaningful comparisons of the identified sets. When the two data sources are drawn from the same population (i.e., covariates are balanced), this assumption is unnecessary because $F_{X}^{\ast}$ can be identified from the experimental data as $F_{X}^{\ast}(\cdot)=F_{X|G}^{\ast}(\cdot\mid\mathrm{exp})$. In other cases, covariate information for the entire population is often available from external sources such as demographic datasets. Moreover, when the parameter of interest is the covariate-conditional effect $\theta_{o,x}$ or the effect $\theta_{o,\mathrm{exp}}$ defined for the experimental population, knowledge of $F_{X}^{\ast}$ is not required.
@@
This set consists of all parameter values that are attainable under some distribution in $\mathcal{F}_{\mathrm{exp}}^{\ast}$. Any parameter value outside $\Theta_{I}$ is incompatible with the experimental data or the maintained assumptions.
We next consider the case in which both experimental and observational data are available. We first define the identified set for the joint CDF $F^{\ast}$ under the combined data, analogous to $\mathcal{F}_{\mathrm{exp}}^{\ast}$, as follows:
@@
where $F_{YDGX}^{\ast}$ is the joint distribution of the observables $(Y,D,G,X)$ in the combined data. This set consists of all CDFs that satisfy the maintained assumptions and are consistent with the observed distribution $F_{YDGX}^{\ast}$. By construction, $\mathcal{F}^{\ast}\subseteq\mathcal{F}_{\mathrm{exp}}^{\ast}$, since the condition $F_{YDGX}=F_{YDGX}^{\ast}$ in $\mathcal{F}^{\ast}$ implies both $F_{YDX|G}(\cdot,\cdot,\cdot\mid\mathrm{exp})=F_{YDX|G}^{\ast}(\cdot,\cdot,\cdot\mid\mathrm{exp})$ and $F_{X}=F_{X}^{\ast}$.
@@
This set consists of all parameter values that are attainable under some CDF in $\mathcal{F}^{\ast}$. Any value outside $\Theta_{IC}$ contradicts the maintained assumptions or the observed combined data.
Since $\mathcal{F}^{\ast}\subseteq\mathcal{F}_{\mathrm{exp}}^{\ast}$, it follows that $\Theta_{IC}\subseteq\Theta_{I}$, meaning that the identified set under the combined data is no larger than that under the experimental data alone. Our primary interest, however, lies in whether the strict inclusion $\Theta_{IC}\subset\Theta_{I}$ holds. This corresponds to asking whether combining experimental and observational data strictly tightens the identified set for $\theta_{o}$. This question is nontrivial because a strict inclusion $\mathcal{F}^{\ast}\subset\mathcal{F}_{\mathrm{exp}}^{\ast}$ at the level of CDFs does not necessarily imply a strict inclusion $\Theta_{IC}\subset\Theta_{I}$ at the level of the parameter. We examine this question in the following sections.
@@ 3 Characterization of the Identified Sets
To investigate whether the strict inclusion $\Theta_{IC}\subset\Theta_{I}$ holds, the definitions of $\Theta_{I}$ and $\Theta_{IC}$ in equations (2) and (3) are too abstract to provide direct insight. We therefore seek to characterize these identified sets in a more interpretable form. To this end, we employ the concept of bivariate copulas (Sklar, 1959), which serves as a central tool for the analysis in this and subsequent sections.
@@ 3.1 Characterization of $\Theta_{I}$ for Experimental Data
We begin with the case of using only experimental data. Under randomized treatment assignment (Assumption 2.3) and external validity (Assumption 2.4), the marginal distributions of the potential outcomes conditional on covariates, $(F_{Y_{1}|X}^{\ast},F_{Y_{0}|X}^{\ast})$, are identified as $F_{Y_{d}|X}^{\ast}(\cdot|x)=F_{Y|DGX}^{\ast}(\cdot|d,\mathrm{exp},x)$ for d=0,1d=0,1.444This follows from the equalities $F_{Y_{d}|X}^{\ast}(\cdot|x)=F_{Y_{d}|GX}^{\ast}(\cdot|\mathrm{exp},x)=F_{Y|DGX}^{\ast}(\cdot|d,\mathrm{exp},x)$, where the first and second equalities follow from Assumptions 2.4 and 2.1, respectively. Since $F_{X}^{\ast}$ is assumed to be known, we can identify the joint distribution of each potential outcome and the covariates, $(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast})$.
Let $\mathcal{C}$ denote the class of all bivariate copula functions. For each $x\in\mathcal{X}$, let $C^{\ast}(\cdot,\cdot|x)\in\mathcal{C}$ denote the true conditional copula given $x$, which reproduces the true conditional joint distribution $F_{Y_{1}Y_{0}|X}^{\ast}(y_{1},y_{0}|x)$ from the marginals $(F_{Y_{1}|X}^{\ast}(y_{1}|x),F_{Y_{0}|X}^{\ast}(y_{0}|x))$ as follows:
@@
where the existence of such a copula function is guaranteed by Sklar’s theorem (e.g., Nelsen, 2006, Theorem 2.3.3). [^5] Sklar’s theorem (Sklar, 1959) states that any joint distribution $F_{Y_{1}Y_{0}}$ can be expressed as a copula function of its marginals: $F_{Y_{1}Y_{0}}(y_{1},y_{0})=C(F_{Y_{1}}(y_{1}),F_{Y_{0}}(y_{0}))$ for some copula function $C$. Conversely, given any marginal distributions $F_{Y_{1}}$ and $F_{Y_{0}}$, the function $C(F_{Y_{1}}(y_{1}),F_{Y_{0}}(y_{0}))$, for any copula $C$, defines a valid bivariate distribution with those marginals. Using the true conditional copula $C^{\ast}(\cdot,\cdot|x)$, the parameter $\theta_{o}$ can be expressed as
@@
The true copula $C^{\ast}(\cdot,\cdot|x)$ is unknown. However, by replacing $C^{\ast}(\cdot,\cdot|x)$ with all possible copula functions $C(\cdot,\cdot|x)\in\mathcal{C}$, we obtain the identified set for $\theta_{o}$ based on the knowledge of $(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast})$ as
@@
By Sklar’s theorem, the collection $\bigl\{C\bigl(F_{Y_{1}|X}^{\ast}(\cdot|x),F_{Y_{0}|X}^{\ast}(\cdot|x)\big|x\bigr):C(\cdot,\cdot|x)\in\mathcal{C}\bigr\}$ coincides with the set of all conditional joint CDFs of $(Y_{1},Y_{0})$ given $X=x$ that share the marginals $\bigl(F_{Y_{1}|X}^{\ast}(\cdot|x),F_{Y_{0}|X}^{\ast}(\cdot|x)\bigr)$.
The following proposition formally establishes that the identified set $\widetilde{\Theta}_{I}$, based on the distributions $(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast})$, coincides with the identified set $\Theta_{I}$ under the experimental data.
@@
The identified set $\Theta_{I}$ under the experimental data is equivalent to the identified set $\widetilde{\Theta}_{I}$ based on the distributions $(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast})$; that is, $\Theta_{I}=\widetilde{\Theta}_{I}$.
This result formally confirms that $\Theta_{I}$ is fully characterized by the marginal distributions $(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast})$. [^6] While many studies refer to $\widetilde{\Theta}_{I}$ as the identified set under experimental data, Proposition 3.1 provides the formal justification for this equivalence.